OpenStudy (anonymous):

Solve the following system of equations: −2x + y = 1 −4x + y = −1

OpenStudy (anonymous):

By substitution or elimination. EQ#1 times -1 2x - y = -1 Add this vertically to EQ#2 -4x + y = -1 ---------------- -2x = -2 x = 1 Now plug this into either original equation to find y. -2(1) + y = 1 -2 + y = 1 y = 3 Solution (1,3)